The sum of two numbers is 18. Twice the smaller number decreased by three equals the larger number. What are the two numbers?
Answer by ankor@dixie-net.com(22740)

The sum of two numbers is 18.
a + b = 18
Twice the smaller number decreased by three equals the larger number.
a = 2b - 3
What are the two numbers?
:
In the first equation, replace a with(2b-3)
2b-3 + b = 18
3b = 18 + 3
3b = 21
b = 7
:
:
You can find a